If $16,000 is invested in an account for 15 years. Calculate the total interest earned at the end of 15 years if the interest is:

(a) 5% simple interest: $

(b) 5% compounded annually: $

(c) 5% compounded quarterly: $

(d) 5% compounded monthly: $

Round your answers to the nearest cent.
Answer #1

a) Simple Interest = ptr = 16000 * 15 * 5/100 = 12000

b) A = P(1 + r)t

        = 16000(1 + 5/100)15

       = 16000(105/100)15

       = 33262.85

CI = A - P

     = 33262.85 - 16000

     = 17262.85

C) A = P(1 + r/n)nt

        = 16000(1 + 5/400)4 * 15

       = 16000(405/400)60

       = 33714.902

CI = A - P

     = 33714.902 - 16000

     = 17714.902

d) A = P(1 + r/n)nt

        = 16000(1 + 5/1200)12 * 15

       = 16000(1205/1200)180

       = 33819.263

CI = A - P

     = 33819.263 - 16000

     = 17819.263